QUESTION 4 Collins Food Group (CFG) owns both Sizzler and KFC in Australia. Due to Sizzlers continuing poor performance, CFG is deciding whether or not to close the Sizzler restaurants and re-invest into KFC to continue its growth. CFGs every three months revenues are $458,228 while Sizzler remains open, and they estimate revenues will increase to $513,139 every three months if the divestment of Sizzler goes ahead. Running costs will remain the divestment of sizzler would be $97,567 over the time frame of one year What will be CFGs estimated accounting profits for the first three months after the divestment goes ahead? Answer number to the nearest whole number and include a negative sign if it is an accounting loss (with no decimal places, S sign, spaces or commas).
Answer #1

Accounting profits = total revenue (TR) - Total cost (TC)

Total cost includes only explicit cost & not implicit cost

So TR from disinvestment for first 3 months = 513,139

Now Total explicit cost = running cost + disinvestment costs ( for three months)

= 80461 + (97567/12)*3

Since disinvestment costs are for one year, so for one month , 97567/12 = 8130.5834

So total cost = 80461 + 24391.75 = 104,852.75

Thus accounting profits = 513,139- 104,852.75

= 408,286.25 $